# Evolution of binaries containing a hot subdwarf and a white dwarf to double white dwarfs, and double detonation supernovae with hypervelocity runaway stars

### v1

@@ 2.3 Conditions for ignition
We evolve our binaries either until He ignition occurs in the accretor or when the non-interacting hot subdwarf evolves into a WD. The thermonuclear explosion of a C/O WD is a multidimensional phenomenon (Pakmor et al., 2024). However, our simulations use hydrostatic 1D MESA calculations, which do not fully capture the multidimensional nature of the explosion or the nature of nuclear burning processes. Therefore, for the binaries in which He ignition occurs in the accretor, we post-process to classify the detonation events. We use an approximate criterion to classify outcomes as either “detonation” or “non-detonation”. In this classification scheme, we assume all “detonations” trigger a second detonation in the core of the WD, leading to a double detonation supernova (Fink et al., 2010; Boos et al., 2021; Gronow et al., 2021). Conversely, “non-detonations” result in a He nova, where the He shell is expelled but the WD itself remains intact.
From the attempts to understand the distinguishing criteria between detonation and deflagration, previous works have shown that this classification is correlated to various factors such as ignition temperature-dependent critical density (Woosley & Kasen, 2011), and density of the ignition region (Woosley & Weaver, 1994). From Woosley & Weaver (1994), the He flashes in the degenerate surface of the WD heavily depend on the density of the ignition region. They found that for the systems undergoing detonation, the minimum density of the He shell was $\rho=6.8\times 10^{5}\,\text{g\,cm}^{-3}$. Furthermore, most of the systems that experienced detonations were characterized by densities above $\rho=10^{6}\,\text{g\,cm}^{-3}$. In this work, we adopt a critical density $\rho_{c}=10^{6}\,\text{g\,cm}^{-3}$. We classify all systems that undergo He ignition at densities below this critical as non-detonations (He novae), and those igniting above this density as detonations leading to thermonuclear supernovae.
@@ 3 Description of Example systems
In this section, we discuss examples of the three potential outcomes for hot subdwarf + WD binaries as double detonation supernovae, He novae, and double white dwarfs. Fig. 1 shows a schematic diagram for three different fates. Double detonation supernovae and He novae result from systems that undergo mass transfer and later experience thermonuclear instability in the He layer on the surface of the WD. We differentiate a double detonation supernova from He nova based on the He layer’s critical density (see Sec. 2.3). Double white dwarf systems result from binaries that either do not come into contact or involve episodes of accretion that fail to ignite the He layers of the WD.
@@ 3.1 Supernova
We first examine a system that evolves into a double detonation supernova. For this example, we initiate a hot subdwarf ($0.5\,\mathrm{M_{\odot}}$) and a WD ($0.8\,\mathrm{M_{\odot}}$) binary system with an orbital period of 0.84 hours, assuming the hot subdwarf has just started core He burning. Figure 2 illustrates the evolution of key system properties: (a) the mass of the donor and the accretor, (b) the orbital period of the binary, (c) the radius of the donor, (d) the mass transfer/accretion rate, (e) the surface mass fractions of H and He in the donor, (f) the He mass in the accretor, and (g) the luminosity of the donor, accretor, and the accretion process which we estimate as accretion luminosity
@@
where $G$ is the gravitational constant, $M_{\text{WD}}$ is the mass of the WD, $\dot{M}$ is the mass transfer rate, and $R_{\text{WD}}$ is the radius of the WD.
@@
Figure 6: Mapping of outcomes (supernova, He nova, and double white dwarf) across the initial parameter space. The x-axis represents the initial orbital period, and the y-axis represents the initial donor mass ($\mathrm{M_{sdO/B}}$). The 8 sub-panels describe different white dwarf masses, as labeled, that increase from the top left to the bottom right. Red points denote double detonation supernovae, orange points indicate He novae and grey points represent double white dwarfs. The stars highlight observed hot subdwarf + WD binary systems (see Table 2), with arrows indicating that initial orbital periods must be longer than the observed values. The lower mass donors ($<0.4\,\mathrm{M_{\odot}}$) only lead to double detonation supernovae with higher mass accretors ($>0.9\,\mathrm{M_{\odot}}$). For details, see Sec. 4.
@@
For the initial 11 million years, gravitational wave radiation shortens the binary orbital period until the onset of mass transfer at $\sim 0.55$ hours. The hot subdwarf has a convective He burning core and an H-rich envelope. Mass transfer in typical hot subdwarf + WD binaries occurs in two phases: envelope mass transfer and He core mass transfer (Bauer & Kupfer, 2021). Initially, the donor transfers its H-rich envelope for about 2 million years at a relatively slow mass-transfer rate of about $\sim 10^{-9}\,\mathrm{M_{\odot}yr^{-1}}$. After the envelope is stripped off, the donor transitions to transferring He for the next 7 million years, with the mass-transfer rate increasing to $\sim 10^{-8}\,\mathrm{M_{\odot}yr^{-1}}$(see panel d of Fig. 2), accumulating a dense He layer ($\sim 0.16\,\mathrm{M_{\odot}}$) on the surface of the accretor. Throughout the accretion process, the rates are too low for stable He burning and growth of the C/O core, instead He accumulates as the He envelope.
@@
In the left panel of Fig. 3, we see the time evolution of the accretor’s temperature-density ($T-\rho$) profile. At the start, the temperature decreases monotonically from the center ($\rho\sim 10^{7}\,\mathrm{g\,cm}^{-3}$) to the surface. After He accretion starts at about $14\,\mathrm{Myr}$, a temperature inversion occurs due to compressional heating from the accretion (e.g., Nomoto 1982a; Townsley & Bildsten 2004). As the accretor continues to accrete He at a steady rate of $\sim 10^{-8}\,\mathrm{M_{\odot}yr^{-1}}$, the combination of compressional heating and temperature diffusion continues to compress and heat the underlying layers, shifting the maximum temperature to deeper He layers with densities $\rho\sim 2\times 10^{6}\,\mathrm{g\,cm}^{-3}$. This eventually leads to thermonuclear instability in the degenerate He layers, resulting in the runaway fusion of He via the NCO chain (marked as He ignition in Fig. 3). We stop our simulation at this point.
@@
To understand the dominant process that contributes to reaching the He ignition conditions, we use the analytic expressions described below. We can estimate the adiabatic compression, at constant entropy, using the definition of the third adiabatic index (Hansen et al., 2004)
@@
where $H=P/\rho g$ is the local pressure scale height, $P$ is the pressure, and $g$ is the gravitational acceleration. $D_{\text{th}}=4acT^{3}/3\kappa\rho^{2}c_{P}$ is the coefficient of thermal diffusion, where $a$ is the radiation constant, $c$ is the speed of light, $\kappa$ is the opacity, and $c_{P}$ is the specific heat at constant pressure.
@@
Comparing the timescales for compressional heating, heat diffusion, and neutrino cooling based on the profiles of these quantities from our MESA model, we estimate that heat diffusion is the dominant process in setting the shape of the temperature profile that drives He ignition in the double detonation supernova case described above.
He ignition in this model occurs at densities $\rho\sim 2\times 10^{6}\,\mathrm{g\,cm}^{-3}$ , which is greater than the adopted critical density (see Sec. 2.3). Thus we classify this He ignition as a detonation and we expect that triggers a subsequent core detonation in the WD (Polin et al., 2019), consequently destroying the WD in a thermonuclear supernova. Immediately before the detonation, the WD has a mass of $\sim 0.96\,\mathrm{M_{\odot}}$ with a thick He shell $\sim 0.16\,\mathrm{M_{\odot}}$ and thick He shell likely will lead to a red, faint transient (Kromer et al., 2010; Polin et al., 2019).
@@ 3.2 He nova
The accretor gains mass at a higher rate ($\sim 10^{-7}\,\mathrm{M_{\odot}yr^{-1}}$) compared to the double detonation supernova example, resulting in the transfer of $0.05\,\mathrm{M_{\odot}}$ of He over 1 million years. We note that the mass transfer rates in this scenario are about an order of magnitude larger compared to the supernova example. At these higher rates, the timescale of compressional heating (Eq. 3) is shorter than the time scale of heat diffusion (Eq. 4). Thus, the accretion compresses and heats the less dense layers, eventually triggering a thermonuclear instability. This leads to unstable He burning via the $3\alpha$ process. We stop our simulations at this point.
The right panel of Fig. 3 depicts the evolution of the $T-\rho$ profile of the accretor and marked regions show the He ignition. Since the He ignition occurs at densities $\rho\sim 2.5\times 10^{5}\,\mathrm{g\,cm}^{-3}$, which is less than the critical density (see Sec. 2.3), we expect that this system does not create a detonation when it ignites. While we do not evolve systems beyond the initial thermonuclear instability, previous studies have found various outcomes for the systems undergoing He novae. Some studies (Yungelson, 2008; Nelemans, 2010; Brooks et al., 2015) have found that for systems that survive the initial thermonuclear instability, subsequent evolution is unlikely to lead to a supernova via double detonation. Instead, the continued mass transfer will lead to weaker He shell flashes. However some studies (Hachisu & Kato, 2001; Neunteufel et al., 2016) have found evidence for He-novae being progenitors of Type Ia supernova. Thus, the deterministic fate of these systems can be stated only after full evolution through all the nova outbursts.
@@ 3.3 Double white dwarf
Figure 7: The evolution of selected single hot subdwarf stars is shown in the HR diagram. The annotated red letters “s” and “e” indicate the start and end of the evolution respectively. Different colors represent different masses of hot subdwarf stars, as indicated by the color bar. The constant radius lines illustrate the radius evolution of the hot subdwarf stars during their He shell burning phase. The inset represents a zoomed-in view of the He main sequence phase. Details of the evolution of hot subdwarf as single stars are explained in Sec. 2.1
@@
The system undergoes one episode of accretion with a rate of $\sim 10^{-7}\,\mathrm{M_{\odot}yr^{-1}}$ depositing $\sim 0.04\,\mathrm{M_{\odot}}$ of He on the surface of the accretor. However, the mass transfer stops afterward as the donor runs out of He and evolves into a WD, resulting in a double white dwarf. The accretor (originally a WD) in this double white dwarf has a massive shell due to accretion, unlike those formed through single stellar evolution. We terminate our simulation when the donor’s radius decreases to $0.03\,\mathrm{R_{\odot}}$ and the system is no longer interacting (for more details, see Sec. 2.2). As seen in panel (b) of Fig. 5, the orbital period increases due to mass transfer. However, the magnitude of this increase is not very significant compared to the overall shortening from the initial orbital period (2 hours). The resulting double white dwarf has masses of $\sim 0.66\,\mathrm{M_{\odot}}$ and $\sim 0.89\,\mathrm{M_{\odot}}$, with an orbital period of 1.75 hours. This double white dwarf is expected to merge in 86 million years. Given that one of the WDs has a mass of $\sim 0.9\,\mathrm{M_{\odot}}$, the example system could potentially lead to a Type Ia supernova triggered by the merger (Pakmor et al., 2021).
@@ 4 The fate of binaries across our parameter space
We construct a grid of binaries varying in initial accretor masses, orbital periods, and donor masses. The ranges of donor mass ($\mathrm{M_{hotsubdwarf}}$), accretor mass ($\mathrm{M_{WD}}$), and initial orbital periods ($\mathrm{P_{i}}$) are listed in Table 1 (see also Sec. 2). Fig. 6 presents the main results, showing the fates of the systems as a function of their initial parameters. Each point represents the outcome of a MESA model of a hot subdwarf + WD binary. Eight panels correspond to different initial WD masses. Within each panel, we vary the initial masses for the donor and different initial orbital periods. Systems are color-coded based on their outcomes: red points indicate systems that end in double detonation supernovae, orange points represent systems that lead to He novae, and grey points denote systems that evolve into double white dwarfs. The stars of different colors mark observed systems within our parameter space (see Table 2, Maxted et al. 2000; Geier et al. 2007, 2013; Kupfer et al. 2020b; Pelisoli et al. 2021; Deshmukh et al. 2024).
In the following subsections, we discuss the donor’s impact on the mass transfer rate, how the accretor affects the mass transfer rate. Lastly, we discuss the two interesting outliers.
@@
ZTFJ2055+4651 $56.34785\pm 0.00026$ $0.41\pm 0.04$ $0.68\pm 0.05$ Kupfer et al. (2020a)
PTF J2238+743015.1 $76.341750\pm 0.000001$ $0.383\pm 0.028$ $0.725\pm 0.026$ Kupfer et al. (2022)
HD 265435 $99.09918\pm 0.00029$ $0.63^{+0.13}_{-0.12}$ $1.01\pm 0.15$ Pelisoli et al. (2021)
CD-3011223 70.53 $0.47^{+0.07}_{-0.06}$ $0.74\pm 0.02$ Geier et al. (2013); Deshmukh et al. (2024)
KPD 1930+2752 137 $0.48\pm 0.03$ $1.0\pm 0.03$ Maxted et al. (2000); Geier et al. (2007)

### v2

@@ 2.1 Conditions for ignition
We evolve our binaries either until He ignition occurs in the accretor or when the non-interacting hot subdwarf evolves into a WD. The modeling of white dwarf thermonuclear explosions is inherently a multidimensional problem. (2024A&A...686A.227P). However, our simulations use hydrostatic 1D MESA calculations, which do not fully capture the multidimensional nature of the explosion or the nature of nuclear burning processes. Therefore, for the binaries in which He ignition occurs in the accretor, we post-process to classify the detonation events. We use an approximate criterion to classify outcomes as either “detonation” or “non-detonation.” In this classification scheme, we assume all “detonations” trigger a second detonation in the core of the WD, leading to a double detonation supernova (2010A&A...514A..53F; 2021ApJ...919..126B; 2021A&A...649A.155G). Conversely, “non-detonations” result in a He nova, where the He shell is expelled but the WD itself remains intact.
From the attempts to understand the distinguishing criteria between detonation and deflagration, previous works have shown that this classification is correlated to various factors such as ignition temperature-dependent critical density (2011ApJ...734...38W), and density of the ignition region (1994ApJ...423..371W). From 1994ApJ...423..371W, the He flashes in the degenerate surface of the WD heavily depend on the density of the ignition region. They found that for the systems undergoing detonation, the minimum density of the He shell was $\rho=6.8\times 10^{5}\,\text{g\,cm}^{-3}$. Furthermore, most of the systems that experienced detonations were characterized by densities above $\rho=10^{6}\,\text{g\,cm}^{-3}$. In this work, we adopt a critical density $\rho_{c}=10^{6}\,\text{g\,cm}^{-3}$. We classify all systems that undergo He ignition at densities below this critical as non-detonations (He novae), and those igniting above this density as detonations leading to thermonuclear supernovae.
@@ 3 Description of example systems
In this section, we discuss examples of the three potential outcomes for hot subdwarf + WD binaries as double detonation supernovae, He novae, and double white dwarfs. Fig. 1 shows a schematic diagram for three different fates. Double detonation supernovae and He novae result from systems that undergo mass transfer and later experience thermonuclear instability in the He layer on the surface of the WD. We differentiate a double detonation supernova from He nova based on the He layer’s critical density (see Sect. 2.1). Double white dwarf systems result from binaries that either do not come into contact or involve episodes of accretion that fail to ignite the He layers of the WD.
@@ 3.1 Supernova
We first examine a system that evolves into a double detonation supernova. For this example, we initiate a hot subdwarf ($0.5\,\mathrm{M_{\odot}}$) and a WD ($0.8\,\mathrm{M_{\odot}}$) binary system with an orbital period of 0.84 hours, assuming the hot subdwarf has just started core He burning. Figure 2 illustrates the evolution of key system properties: (a) the mass of the donor and the accretor, (b) the orbital period of the binary, (c) the radius of the donor, (d) the mass transfer rate, (e) the surface mass fractions of H and He in the donor, (f) the He mass in the accretor, and (g) the luminosity of the donor, accretor, and the accretion process which we estimate as accretion luminosity
@@
where $G$ is the gravitational constant, $M_{\text{WD}}$ is the mass of the WD, $\dot{M}$ is the mass transfer rate, and $R_{\text{WD}}$ is the radius of the WD. We estimated $L_{\text{acc}}$ across the entire parameter space, but we did not find a trend that differentiates the outcomes.
@@
Figure 6: Mapping of outcomes (supernova, He nova, and double white dwarf) across the initial parameter space. The x-axis represents the initial orbital period, and the y-axis represents the initial donor mass ($\mathrm{M_{sdO/B}}$). The 8 sub-panels describe different white dwarf masses, as labeled, that increase from the top left to the bottom right. Red points denote double detonation supernovae, orange points indicate He novae and grey points represent double white dwarfs. The stars highlight observed hot subdwarf + WD binary systems (see Table 2), with arrows indicating that initial orbital periods must be longer than the observed values. The lower mass donors ($0.4\,\mathrm{M_{\odot}}$) only lead to double detonation supernovae with higher mass accretors ($0.9\,\mathrm{M_{\odot}}$). For details, see Sect. 4.
For the initial 11 million years, gravitational wave radiation shortens the binary orbital period until the onset of mass transfer at $\sim 0.55$ hours. In MESA, this inspiral is calculated through the angular momentum loss from gravitational waves, as given by Eq. 4 of 2015ApJS..220...15P (see also 1964PhRv..136.1224P):
$\dot{J}_{\rm gr}=-\frac{32}{5}\frac{G^{7/3}}{c^{5}}\left(\frac{2\pi}{P_{\rm orb}}\right)^{7/3}\frac{(M_{1}M_{2})^{2}}{(M_{1}+M_{2})^{2/3}},$ (2)
@@
where $M_{1}$ and $M_{2}$ are the masses of the two stars, $P_{\rm orb}$ is the orbital period, and $c$ is the speed of light. The hot subdwarf has a convective He burning core and an H-rich envelope. Mass transfer in typical hot subdwarf + WD binaries occurs in two phases: envelope mass transfer and He core mass transfer (2021ApJ...922..245B). Initially, the donor transfers its H-rich envelope for about 2 million years at a relatively slow mass-transfer rate of about $\sim 10^{-9}\,\mathrm{M_{\odot}yr^{-1}}$. After the envelope is stripped off, the donor transitions to transferring He for the next 7 million years, with the mass-transfer rate increasing to $\sim 10^{-8}\,\mathrm{M_{\odot}yr^{-1}}$(see panel d of Fig. 2), accumulating a dense He layer ($\sim 0.16\,\mathrm{M_{\odot}}$) on the surface of the accretor. Throughout the accretion process, the rates are too low for stable He burning and growth of the C/O core, instead He accumulates as the He envelope.
@@
In the left panel of Fig. 3, we see the time evolution of the accretor’s temperature-density ($T-\rho$) profile. At the start, the temperature decreases monotonically from the center ($\rho\sim 10^{7}\,\mathrm{g\,cm}^{-3}$) to the surface. After He accretion starts at about $14\,\mathrm{Myr}$, a temperature inversion occurs due to compressional heating from the accretion (e.g., 1982ApJ...253..798N; 2004ApJ...600..390T). As the accretor continues to accrete He at a steady rate of $\sim 10^{-8}\,\mathrm{M_{\odot}yr^{-1}}$, the combination of compressional heating and temperature diffusion continues to compress and heat the underlying layers, shifting the maximum temperature to deeper He layers with densities $\rho\sim 2\times 10^{6}\,\mathrm{g\,cm}^{-3}$. This eventually leads to thermonuclear instability in the degenerate He layers, resulting in the runaway fusion of He via the NCO chain (marked as He ignition in Fig. 3). We stop our simulation at this point.
@@
To understand the dominant process that contributes to reaching the He ignition conditions, we use the analytic expressions described below. We can estimate the adiabatic compression, at constant entropy, using the definition of the third adiabatic index (hansen2004)
@@
where $H=P/\rho g$ is the local pressure scale height, $P$ is the pressure, and $g$ is the gravitational acceleration. $D_{\text{th}}=4acT^{3}/3\kappa\rho^{2}c_{P}$ is the coefficient of thermal diffusion, where $a$ is the radiation constant, $\kappa$ is the opacity, and $c_{P}$ is the specific heat at constant pressure.
@@
For this system, we estimate the timescales for compressional heating, heat diffusion, and neutrino cooling to be approximately $\sim 10^{7}$ years, $\sim 6\times 10^{5}$ years, and $\sim 5\times 10^{6}$ years, respectively. Based on the profiles of these quantities from our MESA model, we conclude that heat diffusion is the dominant process in shaping the temperature profile that drives He ignition in the double detonation supernova case described above.
He ignition in this model occurs at densities $\rho\sim 2\times 10^{6}\,\mathrm{g\,cm}^{-3}$ , which is greater than the adopted critical density (see Sect. 2.1). Thus we classify this He ignition as a detonation and we expect that triggers a subsequent core detonation in the WD (2019ApJ...873...84P), consequently destroying the WD in a thermonuclear supernova. Immediately before the detonation, the WD has a mass of $\sim 0.96\,\mathrm{M_{\odot}}$ with a thick He shell $\sim 0.16\,\mathrm{M_{\odot}}$ and thick He shell likely will lead to a red, faint transient (2010ApJ...719.1067K; 2019ApJ...873...84P).
@@ 3.2 He nova
The accretor gains mass at a higher rate ($\sim 10^{-7}\,\mathrm{M_{\odot}yr^{-1}}$) compared to the double detonation supernova example, resulting in the transfer of $0.05\,\mathrm{M_{\odot}}$ of He over 1 million years. We note that the mass transfer rates in this scenario are about an order of magnitude larger compared to the supernova example. At these higher accretion rates, the timescale for compressional heating is estimated to be $\sim 124$ years (Eq. 4), which is shorter than the estimated timescale for heat diffusion, $\sim 0.2\times 10^{5}$ years (Eq. 5). Thus, the accretion compresses and heats the less dense layers, eventually triggering a thermonuclear instability. This leads to unstable He burning via the $3\alpha$ process. We stop our simulations at this point.
The right panel of Fig. 3 depicts the evolution of the $T-\rho$ profile of the accretor and marked regions show the He ignition. Since the He ignition occurs at densities $\rho\sim 2.5\times 10^{5}\,\mathrm{g\,cm}^{-3}$, which is less than the critical density (see Sect. 2.1), we expect that this system does not create a detonation when it ignites. While we do not evolve systems beyond the initial thermonuclear instability, previous studies have found various outcomes for the systems undergoing He novae. Some studies (2008AstL...34..620Y; 2010Ap&SS.329...25N; 2015ApJ...807...74B) have found that for systems that survive the initial thermonuclear instability, subsequent evolution is unlikely to lead to a supernova via double detonation. Instead, the continued mass transfer will lead to weaker He shell flashes. However, some studies (2001ApJ...558..323H; 2016A&A...589A..43N) have found evidence for He-novae being progenitors of Type Ia supernova. Thus, the deterministic fate of these systems can be stated only after full evolution through all the nova outbursts.
@@ 3.3 Double white dwarf
Figure 7: Evolution of selected single hot subdwarf stars is shown in the HR diagram. The annotated red letters “s” and “e” indicate the start and end of the evolution respectively. Different colors represent different masses of hot subdwarf stars, as indicated by the color bar. The constant radius lines illustrate the radius evolution of the hot subdwarf stars during their He shell burning phase. The inset represents a zoomed-in view of the He main sequence phase. Details of the evolution of hot subdwarf as single stars are explained in Sect. 2
@@
The system undergoes one episode of accretion with a rate of $\sim 10^{-7}\,\mathrm{M_{\odot}yr^{-1}}$ depositing $\sim 0.04\,\mathrm{M_{\odot}}$ of He on the surface of the accretor. However, the mass transfer stops afterward as the donor runs out of He and evolves into a WD, resulting in a double white dwarf. The accretor (originally a WD) in this double white dwarf has a massive shell due to accretion, unlike those formed through single stellar evolution. We terminate our simulation when the donor’s radius decreases to $0.03\,\mathrm{R_{\odot}}$ and the system is no longer interacting (for more details, see Sect. 2). As seen in panel (b) of Fig. 5, the orbital period increases due to mass transfer. However, the magnitude of this increase is not very significant compared to the overall shortening from the initial orbital period (2 hours). The resulting double white dwarf has masses of $\sim 0.66\,\mathrm{M_{\odot}}$ and $\sim 0.89\,\mathrm{M_{\odot}}$, with an orbital period of 1.75 hours. This double white dwarf is expected to merge in 86 million years. Given that one of the WDs has a mass of $\sim 0.9\,\mathrm{M_{\odot}}$, the example system could potentially lead to a Type Ia supernova triggered by the merger (2021MNRAS.503.4734P).
@@ 4 The fate of binaries across our parameter space
We construct a grid of binaries varying in initial accretor masses, orbital periods, and donor masses. The ranges of donor mass ($\mathrm{M_{hotsubdwarf}}$), accretor mass ($\mathrm{M_{WD}}$), and initial orbital periods ($\mathrm{P_{i}}$) are listed in Table 1 (see also Sect. 2). Fig. 6 presents the main results, showing the fates of the systems as a function of their initial parameters. Each point represents the outcome of a MESA model of a hot subdwarf + WD binary. Eight panels correspond to different initial WD masses. Within each panel, we vary the initial masses for the donor and different initial orbital periods. Systems are color-coded based on their outcomes: red points indicate systems that end in double detonation supernovae, orange points represent systems that lead to He novae, and grey points denote systems that evolve into double white dwarfs. The stars of different colors mark observed systems within our parameter space (see Table 2, 2000MNRAS.317L..41M; Geier2007; Geier2013; 2020ApJ...891...45K; 2021NatAs...5.1052P; 2023MNRAS.525..183S; 2024MNRAS.527.2072D). In the following subsections, we discuss the donor’s impact on the mass transfer rate, how the accretor affects the mass transfer rate. Lastly, we discuss the two interesting outliers.
@@
ZTFJ2055+4651 $56.34785\pm 0.00026$ $0.41\pm 0.04$ $0.68\pm 0.05$ 2020ApJ...898L..25K
PTF J2238+743015.1 $76.341750\pm 0.000001$ $0.383\pm 0.028$ $0.725\pm 0.026$ 2022ApJ...925L..12K
HD 265435 $99.09918\pm 0.00029$ $0.63^{+0.13}_{-0.12}$ $1.01\pm 0.15$ 2021NatAs...5.1052P
CD-3011223 70.53 $0.47^{+0.07}_{-0.06}$ $0.74\pm 0.02$ Geier2013; 2024MNRAS.527.2072D
KPD 1930+2752 137 $0.48\pm 0.03$ $1.0\pm 0.03$ 2000MNRAS.317L..41M; Geier2007
Ton S 415 $84.646\pm 0.0004$ $0.33\pm 0.09$ $0.47\pm 0.24$ 2023MNRAS.525..183S